OpenStudy (anonymous):

write an equation of the line that passes through the given points. (3,-3), (4,1)?

OpenStudy (hoblos):

slope= 4 y-1= 4(x-4) y=4x-15

OpenStudy (anonymous):

slope = 1--3 / 4-3 = 4 using y-y1 = m(x-x1) y - 1 = 4(x-4) y = 4x -15

OpenStudy (anonymous):

thanks